### v1

# When Large Language Models Meet Vector Databases: A Survey

Abstract
The recent burst in Large Language Models has opened new frontiers in human-like text processing and generation. However, alongside their remarkable growth, Large Language Models have encountered critical challenges including issues of hallucination, bias, real-time knowledge updates, and the high costs of implementation and maintenance in commercial settings. Vector Databases, another increasingly popular tool, offer potential solutions to these challenges. These databases are adept at handling high-dimensional data and are crucial for tasks such as efficient information retrieval and semantic search. By integrating with Large Language Models, they significantly enhance AI systems’ ability to manage and utilize diverse data more effectively. This survey paper provides an in-depth and unique analysis of the intersection between Large Language Models and Vector Databases.
1 Introduction
The concept of Artificial Intelligence (AI) was envisioned decades ago, but it is not until ChatGPT unlocked the power of interactive smart AI assistants that people are offered with the unprecedentedly tangible experience of talking to a machine program that ”seems” more knowledgeable than an ordinary person. The phenomenal heat of ChatGPT inherits from the success of Large Language Models (LLMs), one major element of AI. LLMs, exemplified by systems like GPT Brown and et al. (2020); Achiam and et al. (2023), BERT Devlin et al. (2018), and Llama Touvron and et al. (2023), can process, understand, and generate human-like text. Thanks to the power of pre-training which involves training a language model on a massive corpus of text data, LLMs can capture the complexities of human languages, including context, idioms, and even cultural references. Their applications range from simple text completion to complex tasks like translation, understanding, and generation, making them invaluable assets in both commercial and research domains.
However, despite the impressive capabilities of LLMs, they are still under the shadows of doubt in many aspects. One major shortcoming is the problem of hallucination, where LLMs generate plausible but factually incorrect or faithfully nonsensical information Huang et al. (2023). Causes behind this problem include 1. Lack of domain knowledge. The fact that LLMs are primarily trained on public datasets Penedo et al. (2023) inevitably leads to a limited ability to answer domain-specific questions that are out of the scope of their internal knowledge. 2. Real-time knowledge updates. Even if the questions are within the learning corpus of LLMs, their answers may still exhibit limitations because the internal knowledge may be outdated when the outside world is dynamic and keeps changing Onoe et al. (2022). 3. Biases. The datasets used to train LLMs are large which may introduce systematic errors Bender et al. (2021). And essentially every dataset can be questioned with biases issues, including imitative falsehoods Bender et al. (2021), duplicating biases Lee et al. (2022), and social biases Venkit et al. (2023). Moreover, a disadvantage of incorporating LLMs commercially is the expensive cost of maintenance. For an average business entity, applying LLMs for business use is barely feasible. It is almost impossible for a non-tech company to customize and train a GPT model of its own because they don’t have the resources and talents to conduct such big of a project Musser (2023), while frequent API calls to third-party LLM providers like OpenAI can be extremely expensive, not to mention there is a very limited number of such providers in certain areas. Additionally, the oblivion problem of LLMs has been of controversial because LLMs are proven to tend to forget.
Structure Encoder Only Encoder-Decoder Decoder Only
LLMs BERT Devlin et al. (2018), RoBERTa Liu et al. (2019), XLM Lample and Conneau (2019), ALBERT Lan and et al. (2019), ELECTRA Clark and et al. (2020), DeBERTa He and et al. (2020) T5 Raffel et al. (2020), BART Lewis and et al. (2019), mT5 Xue and et al. (2020), M2M-100 Fan and et al. (2021), BigBird Zaheer and et al. (2020), ChatGLM Zeng and et al. (2022) GPT-2 Radford and et al. (2019), GPT-3 Brown and et al. (2020), OPT Zhang and et al. (2022), PaLM Chowdhery and et al. (2023), BLOOM Workshop and et al. (2022), MT-NLG Smith and et al. (2022), GLaM Du and et al. (2022), Gopher Rae and et al. (2021), chinchilla Hoffmann and et al. (2022), LaMDA Thoppilan and et al. (2022), LLaMA Touvron and et al. (2023), GPT-4 Achiam and et al. (2023), BloombergGPT Wu and et al. (2023)
Table 1: Summary of LLMs
Another increasingly popular sub-field of AI is AI databases that support vector data storage and its efficient retrieval at scale, also called vector databases (VecDBs). Whatever kind of multi-modal data that LLMs deal with, searching over many vectors is time-consuming, it is recommended by OpenAI and other LLM vendors to use VecDBs, which always provide LLMs and their applications with cost-effective retrievals and scalable data management.
The intersection of Large Language Models and Vector Databases has been and will still be a burgeoning area of study and application, offering exciting possibilities. This synergy allows for the creation of more powerful, efficient, and versatile AI systems, capable of handling a broad spectrum of tasks with enhanced accuracy and speed. It also paves the way for innovative applications in fields like healthcare, finance, education, and entertainment, where AI can deliver tailored solutions and insights.
In the light of lacking papers that introduce LLMs in the view of Vector Databases, this survey aims to picture how Vector Databases can be potential solutions to refine Large Language Models’ known shortcomings in previous works and hope to offer a unique perspective of future directions in the intersection that is fertile of research potentials. This paper develops as below:
  • Section 2 is presented as introduction to the background knowledge of the two major characters of this paper, LLMs and Vector Databases. Section 2.1 offers a brief view of the popular LLMs and accents on their underlying shortcomings, while Section 3.1 depicts the picture of Vector Databases and their unique capabilities in integration with LLMs.
  • In Section 3, we provide a comprehensive summary of how previous research has effectively combined LLMs with Vector Databases. Section 3.2 delves into a detailed blueprint of the Retrieval-Augmented Generation paradigm in which a unique role that Vector Databases play. Section 3.3 and 3.4 showcases other memory-wise benefits that vector databases can bring for LLMs.
  • Section 4 is dedicated to our analysis of the limitations and unexplored areas at the intersection of LLMs and Vector Databases and highlights the potential future research opportunities in this domain.
2 Background
2.1 Large Language Models
Over the last half-decade, we have witnessed the groundbreaking success of Large Language Models (LLMs), marking a significant milestone in the field of Natural Language Processing (NLP). LLMs have revolutionized our views on the approaches to understanding and generating human languages by machines. There are two key factors in the evolution of Large Language Models, the development of Neural Network architectures and the superpower of pre-training models on extensive data.
2.1.1 Development of Language Models
With the advent of neural networks, the field of NLP has undergone a transformative shift, which began with the introduction of Recurrent Neural Networks (RNNs) Zaremba et al. (2014). RNNs provide a way to process sequences of words and capture temporal dependencies in text. And 2 of its famous variants, Gated Recurrent Units (GRUs) Chung et al. (2014) and Long Short-Term Memory networks (LSTMs) Shi et al. (2015) address the limitations of RNNs in handling problems of vanishing and exploding gradients.
The next pivotal milestone for NLP is the widespread adoption of the Transformer architecture Vaswani et al. (2017) which has set a new standard for language models. The multi-head self-attention modules and cross-attention modules in the encoders and decoders enable the model to capture long-range dependencies, parallel processing, and contextual understanding. Furthermore, the model starts to rapidly grow in size in the Transformer era. More importantly, the Transformer represents a paradigm shift in NLP field, and the success of the Transformer architecture becomes the significant backbone of LLMs, upon which the encoder-only, encoder-decoder, and decoder-only models are built.
Encoder-Only Models
These models are designed to analyze and understand input text. They process the input to create representations, in the form of embedding vectors, that capture the nuances and context of the language. Examples of the models in this category include BERT Devlin et al. (2018) and its derivatives, such as RoBERTa Liu et al. (2019). They excel in tasks that require a deep understanding of language context, such as sentiment analysis, named entity recognition, and question answering where the answers are contained within the given text.
Encoder-Decoder Models
This architecture consists of two parts: an encoder that processes the input text and a decoder that generates the output. These models are particularly effective in tasks that involve transforming an input into a different output, such as machine translation or text summarizing. The Transformer model itself, as we have discussed before, is an example of this category, which has brought astonishing breakthroughs in machine translation. Another significant model is T5 Raffel et al. (2020), which frames all language tasks as a text-to-text problem, showcasing the versatility of the encoder-decoder framework.
Decoder-Only Models
Decoder-only architectures are optimized for generating text. These models take an input and then continue to generate language based on that input. GPT-3 Brown and et al. (2020) and its successor, GPT-4 Achiam and et al. (2023), are quintessential examples of this category. They are particularly adept at tasks that require creative and coherent text generation, such as content creation, storytelling, and even code generation.
We show the LLMs that have received notable attention during the development of language models in Table 1.
3 LLM+VectorDB
Figure 1: Sample: A sample RAG framework that uses vector databases.
3.0.1 Power of Pre-training
The scaling laws proposed by OpenAI Kaplan and et al. (2020) highlight a critical trend: the scaling of Pretrained Language Models (PLMs), in terms of both model size and data volume, leads to significant improvement on downstream tasks. This is evidenced by the development of increasingly larger PLMs, such as the 1.76-trillion-parameter GPT-4 Achiam and et al. (2023) and the 540-billion-parameter PaLM Chowdhery and et al. (2023). Unlike their smaller predecessors, like the 330-million-parameter BERT Devlin et al. (2018) and 1.5-billion-parameter GPT-2 Radford and et al. (2019), these large-scale models demonstrate unique behaviors and emergent abilities Wei and et al. (2022) in various tasks like zero-shot learning, which is extremely challenging to the small-scale models. Researchers started to call the large-scale models in the NLP field “Large Language Models” or “LLMs” to denote these significantly scaled PLMs, in order to distinguish the outstanding performance and gain more traction in academia.
The superpower of pre-training lies in its ability to provide models with a general understanding of the languages, which is then tailored through additional training, known as fine-tuning, for specific tasks unlike traditional deep learning methods Dong et al. (2021); Liu and et al. (2024). The advantage is that the model doesn’t start from scratch when learning a new task; it builds upon a rich, pre-existing foundation of language understanding. And this enables what is known as transfer learning Yosinski and et al. (2014). Knowledge gained during this initial phase can be transferred to a wide range of tasks, even those that the model was not explicitly trained on.
3.0.2 Are we there yet? (Challenge faced by pure LLMs)
However, the success of LLMs does not come without challenges. Issues such as data biases Raffel et al. (2020), which can lead to skewed or unfair outcomes Li and et al. (2023), are of significant concern. Ethical considerations, including privacy problems Yao et al. (2024) and the potential for misuse of generative content Ganguli et al. (2022). Training language models to follow instructions with human feedback, like toxic content, also receives quite a bit of debate. Researchers are arguing that the power of LLMs is very superficial, and it is the large pre-training data Schaeffer (2023) that helps the models gain excellent performance on most of the benchmarks. Hallucinations Ji and et al. (2023) are also found to be a major problem with LLMs, casting doubt on the reliability of their outputs. Additionally, the computational resources required for training and running these models are substantial, posing environmental and accessibility questions Strubell and et al. (2019). As the sizes of both LLMs and the training datasets keep getting bigger, the cost of training and inferring is heavily influenced. It is estimated that training the 11-billion parameter version of T5 Raffel et al. (2020) costs over 1.3 million dollars for a single run, whereas one round of training GPT-3 Brown and et al. (2020) of 175 billion parameters using a Tesla V100 cloud instance requires costs 4.6 million dollars. On top of that, training a BERT-based language model Devlin et al. (2018) using 8 V100 GPUs for 36 hours and used a total of 37.3 kWh which is estimated to be more energy-consuming than a gallon of gasoline in terms of $CO_{2}$ emission Dodge and et al. (2022).
3.1 Vector Databases, the V-factor
While LLMs like ChatGPT are relatively new concepts, database management systems (DBMS) have been thoroughly developed and applied in many aspects in the last 60 years of history, well-recognized for their consistent stability and universality for structured data with fixed formats that excel well with computer storage. However, the development and wide application of deep learning models such as convolutional neural networks He and et al. (2016) and transformers Devlin et al. (2018); Su and et al. (2022), enables the embedding of unstructured multi-modal data like images and text mapped into corresponding fixed-length vector representations, which include the high-dimensional semantic features of original data and the semantic similarities are natively represented by distances between vectors, requiring a new type of DBMS that is specifically designed for handling vector data operations, especially vector search and storage.
There are many kinds of databases, as shown in Table 2, whereas vector databases are the only category of databases that natively support diverse unstructured data with efficient storage, indexing, and retrieval. All these operations could be done based on the vector indexes, which are optimally designed collections of vectors deployed together for ANN search. As a result of various blooming applications on the cloud, various data sources are in many different formats and diverse places. Unlike traditional databases that require structured or semi-structured data that must convey a few restrictions and formats, vector databases are purpose-designed for storing the deep learning embedding of various unstructured data that emerged in real-world applications. On the other hand, distinct from traditional DBMS that searches for exact values within databases, vector databases heavily rely on approximate nearest neighbor (ANN) search for vectors, which searches approximate top-k nearest distance neighbors within the high-dimensional base vector data space that do not necessarily require exact matches.
With the unstructured base data embedded into vectors with high dimensionality, calculating k-nearest neighbors of a given query vector data can be expensive, since it requires distance computation to every point in the dataset and maintaining the top-k results. Such operation would direct to a time complexity of $O(dN+N\log{k})$, where $d$ is the dimensionality and $N$ is the number of vectors, for searching top-k results using pair-wise distance calculation and a heap to keep top-k results.
This calls for a more efficient search technique with satisfying accuracy, the ANN Benchmarks  [^1] https://ann-benchmarks.com/ has showcased the great performance gap between brute-force ANN search and index-enhanced ANN search. Since brute force search is both time-consuming and computationally expensive, vector indexing can solve both of these problems, which optimized for ANN search within VDBMS include tree-based Muja and Lowe (2009); Tao et al. (2009), hash-based Andoni and Indyk (2008), Product Quantization (PQ) Jégou et al. (2011), and graph-based methods Malkov and Yashunin (2018). Among all these indexes, the graph-based Hierarchical Navigable Small Worlds (HNSW) provides state-of-the-art performance and capability with great universality and is widely used within most vector database management systems, including the examples we mentioned in Table 3.
Category Supported Data type Examples
Relational Databases (RDBMS) Structured MySQL, PostgreSQL, Oracle
Document Databases Semi-Structured MongoDB, Couchbase
Time-Series Databases Structured InfluxDB, TimescaleDB
Graph Databases Structured Neo4j, Amazon Neptune
Vector Databases (VDBMS) Unstructured Pinecone, Milvus, Jina, Qdrant
Table 2: Different kinds of mainstream business-level databases.
Name Supported Data Supported Query Vector Index
(Version with year) Vec. Dim. Type Filter Multi-Vector Graph IVF
ChormaDB (2022) 1536 Vec. $\checkmark$ $\checkmark$ $\checkmark$
Manu (2022) 32768 Vec. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Milvus (2021) 32768 Vec. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Pinecone (2019) 20000 Vec. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Weaviate (2019) 65535 Vec. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Qdrant (2021) 4096 Vec. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Amazon OpenSearch (v2.9, 2023) 16,000 Ftx. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Elastic Search (v8.0, 2022) 4096 Ftx. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AnalyticDB-V (2020) $\geq$512 Rel. $\checkmark$ $\checkmark$ $\checkmark$
PostgreSQL-pgvector (2021) 2000 Rel. + Ftx. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MongoDB Atlas (v6.0, 2023) 2048 NoSQL $\checkmark$ $\checkmark$
MyScale (2023) 1536 Rel. $\checkmark$ $\checkmark$
Table 3: Comparison of mainstream all-level databases supporting vector data (the version number represents the first version that supports vector search and its release date)
Traditional full-text search engines, such as Elastic Search and Amazon OpenSearch, are based on term frequency metrics like BM25, which has proved practical for many search applications. However, such search techniques require significant investment in time and expertise to tune them to account for the meaning or relevance of the terms searched. On the other hand, vector databases successfully solved the aforementioned problem, thus the emergence of vector databases (as shown in Table 3) has greatly influenced machine learning systems and their multi-modal applications. An intuitive application using vector search is searching images with an image on search engines built on vector databases Wang and et al. (2021) like Google Image Search  [^2] https://cloud.google.com/vertex-ai/docs/vector-search/overview, which uses one input image to find similar images on the internet. While vector databases boast their unique search capability and efficiency for unstructured data, they are just a backend factor for manifold applications. To maximize their abilities, this leads us to an interesting question, since LLMs use embeddings to represent text as vectors, can developers combine vector databases and LLMs to overcome the aforementioned challenges that inherit in LLM applications? The answer is yes.
As a kind of database encapsulated with vector search in vector data that represents real-world information within high dimensionalities, vector databases are well-capable for retrieval applications Asai and et al. (2023) incorporating LLMs because LLM applications are generally read-intensive, not requiring many write-related changes, especially data deletes Pan et al. (2023). On the other hand, vector databases can efficiently manage and warehouse vector data required and generated by LLMs, thus providing a solid data cornerstone for both LLMs and their applications. While LLMs are often limited by domain knowledge that cannot be uploaded or distributed due to security and privacy concerns, domain knowledge is often unstructured data that can easily embedded into vector databases for efficient local retrieval and further integration with generative AIs. Moreover, the computing and storage resources of vector databases are way cheaper than LLMs, since it does require costly GPUs and TPUs, thus achieving a cost-effective way of fast retrieval and durable (non-volatile) storage.
3.2 Retrieval-Augmented Generation (RAG): VecDB as an External Knowledge Base
3.2.1 Development of RAG Paradigm
As the release of ChatGPT casts spotlight on LLMs to the public world, there is an increasing need of using AI chatbots as query and retrieval agents. But simply loading users’ private data as input to LLMs is found to be incompetent in real-world applications. Because LLMs have been constrained by their limited token counts and the high costs associated with training and fine-tuning for every alterations of data, especially pronounced when dealing with personalized or business-specific responses that require real-time data updates. To address such concerns and need, retrieval-augmented generation (RAG) emerges as a novel solution that addresses the challenges faced by LLMs in integrating and processing large and dynamic data in external databases, where vector databases offer a solution by acting as external memory for LLMs. They allow for the segmentation of private data by converting them into vectors and storing these vectors efficiently for quick retrieval process. Integrating with vector databases enables LLMs to access and synthesize enormous amount of data without the need for constant re-training, thereby overcoming their inherent limitations.
The concept of RAG is devised to be a paradigm, and a common workflow of RAG is illustrated in Figure 1. There are essentially 3 main parts to complete one full run of the systems: data storage, retrieval, and generation.
Data storage means establishing reliable external memory like vector databases, and there are detailed recipes for this process Han et al. (2023). It starts with the data preprocessing, during which the original data is collected, cleaned, integrated, transformed, normalized and standardized. The processed data is chunked into smaller segments because of the contextual limit of LLMs. These segments of data are then converted by an embedding model into vectors, the semantic representations of the data, which are stored in a vector database and will be used for vector search in later steps. A well-developed vector database will properly index the data and optimize the retrieval process. The retrieval part starts by a user asking a question in the form of prompts to the same embedding model, which has generated the vector representation of the stored data, and gaining the vector embeddings of the question. The next step of the process is vector searching inside the vector databases The vector search is essentially computing similarity scores among vectors, and the database then identifies and retrieves the data segments that have the highest similarity scores (top K in most RAG systems) compared to the query vector. These retrieved segments are then converted back from their vector format to their original form, which is typically text of documents. In the genertation part, LLMs are involved to generate the final answers. The retrieved documents, along with the user’s question, are incorporated into a specifically chosen and designed prompt template. This template selection is based on the task type and aims to effectively merge the context with the question to form a coherent prompt. The selected LLM is provided with the prompt and generates the final answer.
3.2.2 RAG with Vector Databases
In the prototype of the RAG paradigm, we use such an idea to overcome the hallucination problems of LLMs by providing domain-specific data with accurate instruction, where the vector database serves as an external knowledge base to warehouse domain-specific data, then LLMs can easily handle massive data owned by users.
Even though LLMs such as ChatGPT now have user-specified GPTs for specific uses, with just prompt engineering, their knowledge bases are still limited to the training data provided by OpenAI. However, by using a vector database, users can pre-filter whatever they want it to look at, whereas that is difficult with prompt-engineering GPT assistants.
3.2.3 Expanding Horizons: Multi-modal Integration and Retrieval Innovations in RAG Systems
Numerous studies have been undertaken to enhance the performance and functionality of RAG systems.
Datasets
RAG was originally developed and evaluated on text-based question-answering (QA) tasks. There are numerous datasets and benchmarks available for evaluating the performance of various RAG systems, where different prompt templates need to be designed for different tasks. Therefore, we classified datasets based on the type of tasks and summarized them below.
  • Single-hop QA involves simple questions where the answer can be obtained through a single-step reasoning process, including TriviaQA Joshi and et al. (2017), Natural Questions Kwiatkowski and et al. (2019), WebQuestions Berant and et al. (2013). Multi-hop QA requires multiple intermediate reasoning steps to reach the final answer. For example, for a question like ”In which state did Obama go to high school?”, one needs to first infer which school Obama attended and then determine the location of that school to obtain the final answer. Datasets for multi-hop QA include HotpotQA Yang and et al. (2018), 2WikiMultiHopQA Ho and et al. (2020), MuSiQue Trivedi and et al. (2022), StrategyQA Geva and et al. (2021). In Multiple-choice QA, the model does not need to generate an answer independently but rather select the correct answer from provided options. Multiple-choice QA datasets include MMLU Hendrycks and et al. (2020), SQuAD Rajpurkar and et al. (2016). Open-domain QA involves the task of searching for answers to information-seeking queries within a large pool of knowledge sources. Relevant datasets for open-domain QA include MS MARCO Nguyen and et al. (2016), PopQA Mallen and et al. (2022), CommonsenseQA Talmor and et al. (2018).
  • Fact-checking requires the model to first select relevant sentences from a given document as evidence and then validate statements based on that evidence. Datasets include FEVER Thorne and et al. (2018), FM2 Eisenschlos and et al. (2021).
  • Embedding evaluation comprehensively assess all essential capabilities of text embeddings, including BEIR Thakur and et al. (2021), C-MTEB Xiao and et al. (2023).
Multimodality of RAG
RAG has now evolved to handle a wide range of data types by lending the power of multimodal models.
The impressive achievements of LLMs have inspired significant advancements in vision-and-language research. DALL-E from OpenAI introduced a Transformer-based approach for converting text to images, treating images as sequences of discrete tokens. Subsequent improvements in the text-to-image area Zhang et al. (2023) have been achieved through methods like model scaling, pretraining, and enhanced image quantization models. BLIP-2 Li et al. (2023) uses static image encoders with LLMs for efficient visual language pre-training, facilitating direct image-to-text transformations. Flamingo Alayrac et al. (2022) presented a visual language model for text generation, showcasing remarkable adaptability and leading performance across various vision-and-language tasks. CM3 Aghajanyan et al. (2022) trained a randomly masked model on a large HTML corpus, and showed that the model is capable of generating images and text. FROMAGe Koh et al. (2023) gains robust multimodal capabilities for few-shot learning solely from image-caption pairs, unlike other models that necessitate large-scale, interwoven image-text data from the websites.
To import speech data to RAG systems, Wav2Seq Wu et al. (2022) allows for efficient pre-training without the need for transcriptions, using techniques like k-means clustering and byte-pair encoding to generate pseudo subwords from speech. The Large Language and Speech Model (LLaSM) Shu et al. (2023) is an end-to-end trained, large multi-modal speech-language model equipped with cross-modal conversational skills and proficient in understanding and responding to combined speech-and-language directives. Videos are also made available to certain types of RAG systems. Vid2Seq Yang et al. (2023) enhances language models with specific temporal indicators for predicting event limits and textual descriptions in a single output sequence.
Retrieval Optimizations of RAG
To better harness the knowledge from various types of data, kNN-LMs Khandelwal et al. (2020) explores how incorporating nearest neighbor search into language models can enhance their ability to generalize by effectively leveraging memorized examples. EASE Nishikawa et al. (2022) is distinctive in its use of entities as a strong indicator of text semantics, providing rich training signals for sentence embedding. In-context RALM Ram et al. (2023) proves that with the language model architecture unchanged, simply appending grounding documents to the input will improve the performance. SURGE Kang et al. (2023) enhances dialogue generation by incorporating context-relevant sub-graphs from a knowledge graph. Another work that combines knowledge graphs and LLMs is RET-LLM Modarressi et al. (2023), which is designed to equip large language models with a general write-read memory unit. These studies have focused on retrieval granularity and data structuring levels, with coarse granularity providing more, but less precise, information. Conversely, structured text retrieval offers detailed information at the cost of efficiency.
To utilize both internal knowledge and external resources, SKR Yu et al. (2023) improves LLMs’ ability by enabling them to assess what they know and identify when to seek external information to answer questions more effectively. Selfmem Cheng et al. (2023) enhances retrieval-augmented text generation by creating an unbounded memory pool using the model’s own output and selecting the best output as memory for subsequent rounds. FLARE Jiang et al. (2023) uses predictions of upcoming sentences to anticipate future content and retrieve relevant documents for regenerating sentences, especially when they contain low-confidence tokens. Atlas Izacard et al. (2022) demonstrates impressive performance in tasks like question-answering and fact-checking, outperforming much larger models despite having fewer parameters. Many other works like these also aim to make RAG system more efficient and competent.
3.3 VecDB as a Reliable Memory of GPTs
Oblivion
When using LLM Q&A applications like ChatGPT, LLMs are likely to completely forget the content and information of your previous conversation, even within the same chat tab.
The integration of Large Language Models (LLMs) such as the Generative Pre-trained Transformer (GPT) with vector databases (VecDB) heralds a transformative leap in the field of information retrieval and artificial intelligence. This subsection delves into the compelling synergy between VecDB as a persistent, reliable memory layer and the dynamic, contextually aware computation layer provided by LLMs. In recent works of vector databases that integrate with LLMs, Zhang and et al. (2023) showcases that vector databases’ capability of storing vectors for GPT’s memory.
Vector databases serve as a robust memory layer that addresses one of the intrinsic limitations of LLMs: the static nature of their knowledge. While LLMs excel in generating human-like text based on patterns learned during training, they cannot inherently update their knowledge base dynamically. VecDBs bridge this gap by offering a storage solution that can be continually updated with new information, ensuring that the LLM’s responses are informed by the most current and relevant data available.
The VecDB and LLM combination brings forth a synergy where the LLM provides context and understanding for user queries, while the VecDB offers a precise mechanism for storing and retrieving the relevant vectors. This collaborative approach allows for more accurate, relevant, and efficient responses to complex queries, which would be challenging for either system to address independently.
A VecDB integrated with an LLM facilitates real-time learning and adaptation. As new data is ingested into the VecDB, the LLM can immediately leverage this updated repository to refine its responses. This capability is pivotal for applications requiring up-to-the-minute accuracy, such as financial analysis, news dissemination, and personalized recommendations.
3.4 VecDB as a Cost-effective Semantic Cache
The combination of vector databases and LLM not only facilitates the profound application of LLM with RAG but also provides a new frontier for cost-effective end-to-end LLM applications:
Outrageous API Costs
LLM-based chatbots and agent systems heavily rely on LLM’s output from API vendors, considerable repeated or similar inquiries may lead to outrageous API costs.
API Bandwidth Limitations
Such chatbots and agent systems could also experience bursty inquiries workload that may drown the system’s bandwidth with explosive API calls coming within seconds, leading to the system’s outage and reconfiguration.
One of the primary benefits of integrating vector databases with LLMs is the significant reduction in data operational costs. GPT-Cache, as shown in Figure 2, for instance, stores responses to previously asked queries, and works as a cache before calling LLM APIs. This caching mechanism means that the system doesn’t really need to have API calls to wait for generated responses from scratch every time, reducing the number of costly API calls to the LLM. Moreover, this approach also speeds up the response time, enhancing user experience.
Figure 2: An overview of semantic cache for GPTs that utilizes vector database
Vector databases enhance the LLMs’ ability to retrieve and utilize relevant information by indexing vast amounts of previous Q&A data and mapping them into a vector space, instead of caches in computer systems that require exact hash-match, this allows more precise semantic matching of queries with existing knowledge and results in responses that are not only generated based on the LLM’s training data but also informed by the most relevant and recent information available in the vector database.
4 Discussion: Challenges and Future Work
4.0.1 Are vector searches all you need?
Although vector databases offer a cost-efficient modality for information retrieval within LLM frameworks, their utility in traditional relational database operations remains limited. Specifically, vector-based search methodologies are often not optimized for operations such as post-query filtering, comprehensive full-text searches, and nuanced keyword search mechanisms that are fundamental to conventional database systems. This distinction underscores a potential gap in the functional alignment of vector databases with established data retrieval paradigms, necessitating the development of hybrid search algorithms that can seamlessly integrate vector search with traditional relational database capabilities.
4.0.2 VDBMS with multi-modality
While uni-modal data may lack the prospect of providing informative and more contextually appropriate search results, the multi-modal data and its hybrid processing also present a non-trivial challenge for vector databases’ storage and retrieval Wang and et al. (2023), since integrating and merging multi-modal data require efficient preprocessing with multi-modal encoders, multi-modal storage indexes, but also weight assignment and multi-modal data fusion.
4.0.3 Data preprocessing
While text embedding is considered efficient for processing long text and its retrieval, it comes with a dazzling challenge for building every text-based knowledge database. It is observed that vector retrieval would have fundamentally better performance when precise chunks of text with clear meaning are applied, this raises an interesting question: how to have a proper (probably unified) embedding methodology for every raw text?
On the other hand, dimension reduction for vector data is also considered important due to the curse of dimensionality theory. The current vector data dimension may be not cost-effective, as vectors are large: e.g., a 1536 dimensional vector is about 6 kilobytes, with scaling up to 1 billion, the data size would be 6 terabytes. This requires a more efficient vector data dimension reduction algorithm that matches vector databases and search algorithms.
4.0.4 Knowledge conflict
Conflicting knowledge from the same or different knowledge bases presents a non-trivial challenge for both humans and LLMs to distinguish the correct piece of knowledge. This conflict becomes particularly pronounced when integrating multiple knowledge bases, each potentially carrying its own biases or inaccuracies. Resolving such conflicts requires robust conflict resolution strategies that can assess the reliability of sources and the context of data to determine the most accurate information.
4.0.5 Data management systems for LLMs with multi-tenancy
This challenge is particularly in maintaining the isolation and security of distinct tenants’ data while ensuring efficiency, where isolation is crucial for privacy and security, yet it must be balanced with the need for resource sharing to ensure cost-effectiveness.
5 Conclusion
In this paper, we present a systematic review of recent advances in combinations of LLMs and vector databases. We also introduce recent VecDB+LLMs applications with distinct prototypes that categorize existing works from various perspectives and interdisciplinary studies. Our study also demonstrates both the research and engineering challenges in this fast-growing field and suggests directions for future work.